Subject: APO 3.5m users committee minutes Sept 2014

From: Suzanne Hawley

Submitted: Fri, 10 Oct 2014 20:46:14 -0700 (PDT)

Message number: 1344 (previous: 1343, next: 1345 up: Index)

PLEASE READ: science slides are needed for 3.5m report at ARC Board 
meeting. Send powerpoint slides to your users committee representative 
before October 15.

New ARCSAT schedule (Oct 20-Feb 8) is now published on the APO website.


 		APO 3.5-m Users Committee Phonecon, 9/29/14

Attending:  Suzanne Hawley, Mark Klaene, Scott Anderson, Rachel Kuzio de 
Naray, Mary Beth Kaiser, Rene Walterbos, Jon Holtzman, Michael Strauss, 
Joanne Hughes Clark, John Stocke, Russet McMillan


User feedback and comments from institutional representatives:

The only feedback was from Mary Beth Kaiser who mentioned that some JHU 
users had trouble installing TUI on linux machines.  Jon Holtzman said 
some minor issues were common with linux installs often related to 
accessing python.  Mary Beth was encouraged to contact Russell Owen if 
there are specific questions to address, and to provide feedback regarding 
the online linux installation directions so they can be updated as needed.


Discussion of telescope/instruments report:

Mark presented the telescope and instruments report (below) and mentioned 
the weather as being particularly bad this month.  The monsoon season has 
continued through September with remnants of several hurricanes bringing a 
large amount of rain to the area.  There was a question about the ARCTIC 
schedule.  We are planning for commissioning during summer 2015 and 
possible first shared-risk observations in fall 2015, but will provide 
updates on the schedule as the time gets closer.


 	      3.5-m Telescope and Instruments Highlights
 	                08/16/14 through 09/24/14

0) Overview

Clouds and rain have been the primary issues for much of this period. 
Tropical moisture continues to flow into the region.

Gretchen Van Doren who has been with APO for 24 years is retiring at the 
end of the year. A search for a replacement is in progress however finding 
someone to fill her shoes will likely be impossible.

1) Telescope

The new TCC is working well and aside from 1 item that needs to be tested 
on sky this project has been completed and will be transitioned to 
maintenance mode.

2) Instruments

DIS has been operating normally, but may have some contamination on red 

Agile is operational with no problems. Guiding with new TCC now working.

TripleSpec had no operational problems this period

SPICam was operational during this period with no issues

Echelle inter-order light increased significantly 24 hours after being 
cooled down in August, likely due to condensation/contamination on the 
chip.  Attempts to use the chip heater while keeping the instrument cold 
proved ineffective in improving this. The camera was warmed and held above 
ambient for several days then re-cooled. This improved the inter-order 
light to levels as good or better than pre-shutdown. Condensation was 
observed on ECAM and purge air was routed in to eliminate this.

NIC-FPS continues to operate in shared-risk observing mode due to 
occasional image corruption issue.  NFS less than 9 is operational.

GIFS is operational.  The GSFC folks will be onsite this week to work on 
the instrument.


The instrument development is proceeding on schedule. The CCD has been 
completed and final vendor testing is underway. The dewar assembly and 
initial vacuum testing are complete.  The PO is being issued to Bob Leach 
(ARC Inc) for electronics integration, which should start next week.  We 
have 2 of the 3 quotes back for the optics.


DIS red camera contamination

Similar to the echelle issue above, the red camera on DIS showed evidence 
of contamination due to moisture in the dewar.  We suspect these issues 
are related to the high humidity the past two months.  DIS was taken 
offline for several days and warmed, and now looks better.  John Stocke 
asked about the effect on the data which is most noticeable in the spatial 
direction, but does also affect the focus in the spectral direction.

Science slides for Board annual report

A couple of science slides are needed from each institution for the Board 
report.  They should be powerpoint, with nice pictures if possible and not 
too much technical detail.  Users committee members should gather and send 
to Mark and Suzanne by October 15.  This is an action item.

ARCSAT shutdown and Q4 call for proposals

ARCSAT shutdown to realuminize mirror and work on software issues is 
planned for Sept 29 - Oct 13.  Call for proposals for Q4+ observing (mid 
Oct - early Feb) will go out this week and new schedule should be 
available by Oct 10.

NRC O/IR committee

The NRC has convened an "OIR System Study" committee that is looking at 
the role of US optical and IR ground-based telescopes of medium and large 
aperture in the next decade, particularly in the LSST era.  We have been 
asked to participate in the next committee meeting in mid-October, and 
Rene will be presenting information about the 3.5m telescope 
instrumentation and operations and about the ARC collaboration.  We will 
emphasize time domain and flexible remote operations as usual.  Input 

APOGEE fiber link to 3.5m

SDSS-4 is planning on an APOGEE-S instrument so the current APOGEE 
instrument is likely to stay in the north at APO.  We are therefore 
investigating once again the feasibilty of a fiber link from the 3.5m. 
However since APOGEE will be observing during both bright and dark time, 
it will be difficult to schedule on the 3.5m during SDSS-4 operations (ie 
until 2019-20).  We are interested in hearing from users whether a fiber 
feed to APOGEE is of strong scientific interest and whether we should 
continue to pursue this.  If we do put in fiber to connect the 2.5m and 
3.5m, we can also include the smaller telescopes (ARCSAT, NMSU 1m), so we 
are interested in hearing of any interest in using those telescopes with 
fiber feeds as well.

Action: users committee to gather information from users about scientific 
interest in a fiber feed to APOGEE (2.5m) and smaller telescopes on site.

Initial planning for new spectrograph

With the new imager progressing well, we are starting to consider what the 
next instrument would be.  A new spectrograph to replace DIS is a leading 
candidate, with improvements such as better throughput, higher time 
resolution, and multi-object capability (among others).  We are just 
beginning discussion of this idea, perhaps for a start in 2016.  One 
possibility is to construct a new instrument room next to the arcade, and 
feed a spectrograph from a side port (like Agile) with fibers.  This would 
allow quicker change than mounting on the NA2 port.  Rene mentioned 
possibly making the echelle be fiber fed and freeing up NA1 for a new 
spectrograph.  We will be soliciting input from users in the coming 

Forest service proposal for new site infrastructure

To accomplish the APOGEE fiber feed and/or new instrument room will 
require new site infrastructure, which requires a proposal to the forest 
service. This could take up to a year to get approval, so is a long 
lead-time item.  Mark and APO staff have started work on a proposal while 
we continue discussions about instruments in the meantime.


[open from previous meeting]:


[new from this meeting]:

Action: Users committee representatives to gather science slides for Board 
report and send to Suzanne and Mark by October 15.

Action: Users committee representatives to solicit expressions of 
scientific interest in a fiber feed to APOGEE (2.5m) and smaller 
telescopes and report at next meeting.


Next meeting:  The next Users Committee phonecon will be on Monday 20 
October 2014.  The agenda and other materials will be sent to the 
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1344 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to